# Schur’s Positive-Definite Network: Deep Learning in the SPD cone with structure

### v1

@@ Abstract
Estimating matrices in the symmetric positive-definite (SPD) cone is of interest for many applications ranging from computer vision to graph learning. While there exist various convex optimization-based estimators, they remain limited in expressivity due to their model-based approach. The success of deep learning has thus led many to use neural networks to learn to estimate SPD matrices in a data-driven fashion. For learning structured outputs, one promising strategy involves architectures designed by unrolling iterative algorithms, which potentially benefit from inductive bias properties. However, designing correct unrolled architectures for SPD learning is difficult: they either do not guarantee that their output has all the desired properties, rely on heavy computations, or are overly restrained to specific matrices which hinders their expressivity. In this paper, we propose a novel and generic learning module with guaranteed SPD outputs called SpodNet, that also enables learning a larger class of functions than existing approaches. Notably, it solves the challenging task of learning jointly SPD and sparse matrices. Our experiments demonstrate the versatility of SpodNet layers.
@@ 1 Introduction
The estimation of symmetric positive-definite (SPD) matrices is a major area of research, due to their crucial role in various fields ranging such as optimal transport (Bonet et al., 2023; Ju and Guan, 2022; Yair et al., 2019), graph theory (Lauritzen, 1996), computer vision (Nguyen et al., 2019) or finance (Ledoit and Wolf, 2003). While various statistical estimators, i.e. model-based, have been developed (Ledoit and Wolf, 2004; Banerjee et al., 2008; Cai et al., 2011), a more recent trend is to use learning-based approaches to estimate SPD matrices with neural networks (Gao et al., 2020).
@@
Training neural networks while enforcing non-trivial structural constraints such as positive-definiteness is a difficult task. There have been many efforts in this direction in recent years, often in an ad-hoc manner and each with their own shortcomings (see Section 2 for more details). Building on the seminal work of Gregor and LeCun (2010) in sparse coding, a promising research direction involves designing neural networks architectures from the unrolling of an optimization algorithm (Chen and Pock, 2016; Monga et al., 2021; Chen et al., 2022). In the case of SPD matrices, algorithm unrolling presents several challenges. First, algorithms operating in the SPD cone usually rely on heavy operations such as retractions (Boumal, 2023), SVD or line search (Rolfs et al., 2012). These operations do not integrate well into a neural network architecture, making these algorithms difficult to unroll.
Second, beyond ensuring positive definiteness, it is often desirable to impose additional structure on the matrix to estimate, such as elementwise sparsity. Simultaneously enforcing SPD + sparsity constraints on the output of a neural architecture is extremely difficult (Guillot and Rajaratnam, 2015; Sivalingam, 2015). So far, it is only a hypothetical property that is not satisfied in practice (see Figure 2 and Section 5 for more details).
Contributions: In this paper, we combine the strengths of both model-based and data-driven approaches and we make the following contributions:
  • We introduce a novel framework for building learning modules with a guaranteed SPD-to-SPD mapping for which it is also possible to enforce additional constraints on the output. We refer to it as SpodNet for Schur’s Positive-Definite Network. Our approach is originally inspired by the block-matrix updates perspective of the Graphical Lasso solver (Mazumder and Hastie, 2012) and extended to generic SPD learning. As a particular case, SpodNet layers are able to learn jointly SPD and sparse matrices in a data-driven fashion. To the best of our knowledge, SpodNet-based architectures are the first to offer strict guarantees combining both properties.
@@
  • We illustrate the framework’s versatility through applications in sparse precision matrix estimation. In this scenario, we present limitations of other learning-based approaches and demonstrate how to overcome them with SpodNet layers. Our experiments validate the use of SpodNet for learning jointly SPD-to-SPD and sparsity inducing functions, and show promising results across various performance metrics.
@@
We reserve the bold uppercase for matrices $\bm{\Theta}$, bold lowercase $\bm{\theta}$ for vectors and standard lowercase $\theta$ for scalars. The soft-thresholding function is $\operatorname{ST}_{\gamma}(\cdot)=\operatorname{sign}(\cdot)\max(\lvert\cdot\rvert-\gamma,0)$ for $\gamma\geq 0$; it acts elementwise on vectors or matrices. On matrices, $\lVert\cdot\rVert_{1}$ is the sum of absolute values of the matrix coefficients. The cone of SPD matrices in dimension $p$ is denoted $\mathbb{S}^{p}_{++}$. We use $\bm{\Sigma}$ to designate covariance matrices, $\bm{\Theta}$ for precision matrices, and $\bm{S}$ for empirical covariance matrices.
@@ 2 Related works
Before delving into the details of SpodNet, we briefly review existing approaches for constructing SPD matrices, which can be broadly divided into three categories.
Riemannian learning
Riemannian optimization provides tools to build algorithms whose iterates lie on Riemannian manifolds, such as the SPD manifold (Absil et al., 2008). Many have adopted these tools to design neural architectures that operate on the manifold of SPD matrices (Huang and Van Gool, 2017; Gao et al., 2020, 2022). However, a common and major bottleneck of these models is their dependence on Riemannian operations which are notoriously costly to compute. Furthermore, Riemannian optimization beyond smooth objective functions is still an active area of research (Absil and Hosseini, 2019; Huang and Wei, 2022; Zhou et al., 2023) and it is thus highly non-trivial to design neural architectures to handle additional constraints such as elementwise sparsity.
Projection-based methods
For some applications, it is of interest to project the matrices from a high-dimensional SPD manifold to a lower-dimensional one (Dong et al., 2017; Nguyen et al., 2019). Using linear algebra principles, these approaches design specific layers to guarantee that the final matrix is SPD. However, these solutions are ad-hoc and lack straightforward extensions for general SPD-to-SPD learning with additional structural constraints. Specifically, there is no obvious way to incorporate elementwise sparsity on the matrices without risking breaking their SPD structure. For instance, hard-thresholding the off-diagonal entries of a SPD matrix does not always preserve the SPD property (Guillot and Rajaratnam, 2012) and, more generally, elementwise functions preserving this property are very limited (Guillot and Rajaratnam, 2015).
Unrolled learning
In order to ensure that a network’s output strictly respects some desired properties, another recent and successful direction is to unroll convex optimization algorithms (Chen et al., 2022). This unrolling procedure acts as an “inductive bias” on the architecture (Shrivastava, 2020), naturally guiding the model to explore suitable solutions within the space of imposed properties. In SPD learning, this approach has been exploited by Shrivastava et al. (2019) who unrolled an optimization algorithm to train neural networks to estimate sparse inverses of covariance matrices. Whilst algorithm unrolling is a very powerful approach to learn specific matrices, it is very difficult to extend it to generic SPD learning. We provide more details about the limits of related work in this setting in Section 4.2.
@@ 3 The SpodNet framework
We now introduce the SpodNet layer which is at the core of our contribution. Essentially, it is an SPD-to-SPD mapping parameterized by neural networks. As described in Section 4, SpodNet is inspired by an algorithm used to solve the Graphical Lasso in Mazumder and Hastie (2012).
A SpodNet layer operates by cycling through the $p$ column-row pairs individually and i) updating the corresponding column-row with a neural network, and ii) updating the diagonal element to maintain the SPD constraint (see Figure 1). Interestingly, the neural network at step i) can update the column-row to any value without compromising the SPD guarantee. Consequently, one is free to exploit a spectrum of approaches for those updates depending on other desired structural properties of the output matrix. In particular, we describe in Section 4 a specific implementation of SpodNet that enforces elementwise sparsity for learning sparse precision matrices.
@@
Algorithm 1 One SpodNet layer
Figure 1: Schematic overview of SpodNet layers. Our framework chains updates of column-row pairs and diagonals using neural networks (note: the updated columns and rows are not constant, despite being represented by constant colors). At all times the matrices remain SPD via Schur’s condition.
@@ 3.1 Algorithmic foundations
More formally, when the SpodNet layer operates on the $i$-th column and row of $\bm{\Theta}\in\mathbb{S}_{++}^{p}$ it also updates the $i$-th diagonal coefficient accordingly. The key to preserving the positive-definiteness of the matrix upon changing its $i$-th column and row is an appropriate update of the diagonal entry $\Theta_{ii}$ based on Schur’s condition for positive-definiteness. In the following, a $+$ superscript on a variable (e.g. $\bm{\theta}^{+}$) indicates an update value of this variable (e.g. outputted by a neural network). The following proposition formalizes the SpodNet operations.
@@
(for a generic column $i$, $\bm{\Theta}_{11}$ refers to $\bm{\Theta}$ without its $i$-th row and $i$-th column, $\bm{\theta}_{12}$ is the $i$-th row of $\bm{\Theta}$ without its $i$-th value, and $\theta_{22}$ is $\Theta_{ii}$ as illustrated in Algorithm 1).
Let $f:\mathbb{R}^{p\times p}\rightarrow\mathbb{R}^{p-1}$, $g:\mathbb{R}^{p\times p}\rightarrow\mathbb{R}_{++}$, where $f$ and $g$ are arbitrary functions. Then, updating the $i$-th row and column of $\bm{\Theta}$ as
$\bm{\Theta}^{+}\triangleq\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}^{+}_{12}\triangleq f(\bm{\Theta})\\ \bm{\theta}_{21}^{+}\triangleq f(\bm{\Theta})^{\top}&\theta_{22}^{+}\triangleq g(\bm{\Theta})+\bm{\theta}^{+}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}^{+}_{12}\end{bmatrix}\,,$ (2)
@@
Since $\bm{\Theta}$ is positive-definite, so is its leading principal submatrix $\bm{\Theta}_{11}$. It follows that $\bm{\Theta}^{+}$ is well-defined, and obviously symmetric. Its positive-definiteness then ensues from Schur’s condition for positive-definiteness. Indeed $\bm{\Theta^{+}}$ is SPD when $\bm{\Theta}_{11}\succ 0$ and $\theta_{22}^{+}-\bm{\theta}_{12}^{+\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}=g(\bm{\Theta})>0$ (Zhang, 2006, Theorem 1.12). The latter condition is ensured as the output of $g$ is a positive scalar. ∎
Finally, one SpodNet layer chains $p$ updates of the form (2), sequentially updating all column-row pairs one after the other as summarized in Algorithm 1. A full SpodNet architecture then stacks $K$ SpodNet layers. The overall procedure is schematized in Figure 1 for $K=2$.
The expressivity of SpodNet comes from the simple realization that one is free to plug any arbitrary functions $f$ and $g$ in the updates whilst guaranteeing that $\bm{\Theta}$ remains in the SPD cone. Namely, additional structure such as sparsity can trivially be imposed on $\bm{\Theta}$ through $f$. We next detail the computational cost of fully updating $\bm{\Theta}$ through one SpodNet layer.
3.2 Improving update complexity
The update of Equation 2 a priori requires inverting the matrix $\bm{\Theta}_{11}$ which comes with a prohibitive cost of $\mathcal{O}(p^{3})$ for each column-row update. We can fortunately leverage the column-row structure of the updates to decrease the cost to a mere $\mathcal{O}(p^{2})$, by maintaining the matrix $\bm{W}=\bm{\Theta}^{-1}$ up-to-date along the iterations.
@@ 3.2 Improving update complexity
Let $\bm{W}$ be the inverse of $\bm{\Theta}$, adopting the same block structure $\bm{W}=\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{12}^{\top}&w_{22}\end{bmatrix}$. Then $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$. In addition, if $\bm{\Theta}$ is updated as $\bm{\Theta}^{+}$ following Equation 2, then the update of $\bm{W}$ defined by
$\bm{W}^{+}\triangleq\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}\bm{\theta}_{21}^{+}[{\bm{\Theta}}_{11}]^{-1}}{g(\bm{\Theta})}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}}{g(\bm{\Theta})}\\ \cdot&1/g(\bm{\Theta})\end{bmatrix}\,.$ (3)
@@
By using the Banachiewicz inversion formula on Schur’s complement (Zhang, 2006, Theorem 1.2)
$\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{21}&w_{22}\end{bmatrix}$ $=\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix}^{-1}=\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\\ \\ \cdot&\frac{1}{\theta_{22}-\bm{{\theta}}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\end{bmatrix}\,.$ (4)
Identifying all blocks yields $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$ which can be computed in $\mathcal{O}(p^{2})$ if one has access to $\bm{W}$. This can be achieved using Equation 3, and involves only operations in $\mathcal{O}(p^{2})$. The property $\bm{\Theta}^{+}=[\bm{W}^{+}]^{-1}$ is statisfied by using the same inversion formula. ∎
Based on this result a full update of $\bm{\Theta}$ can be achieved with cost $\mathcal{O}(p^{3})$ ($p$ column-row updates of cost $\mathcal{O}(p^{2})$).
4 Learning sparse precision matrices with SpodNet
In this section, we propose three specific SpodNet-based architectures for learning sparse precision matrices, which naturally arise in the context of Gaussian graphical models (Lauritzen, 1996) that we now briefly recall.
4.1 A model-based formulation: the Graphical Lasso
Consider a dataset of $n$ observed signals $\bm{x}_{1},\ldots,\bm{x}_{n}$ where each $\bm{x}_{i}\in\mathbb{R}^{p}$ follows a certain (centered) distribution with covariance matrix $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$. The problem of precision matrix estimation arises when attempting to identify the conditional dependency graph of the $p$ variables of this dataset. Estimating these dependencies is of interest in many fields such as neuroscience (Rubinov and Sporns, 2010), finance (Ledoit and Wolf, 2003) or genomics (Cai et al., 2013). In Gaussian graphical models, this graph is associated with the so-called precision matrix $\bm{\Theta}=\bm{\Sigma}^{-1}$ which directly encodes the dependencies [^1] In this case $\Theta_{ij}=0$ iff variable $i$ and $j$ are conditionally independent, given the other variables.. The problem of precision matrix estimation involves determining $\bm{\Theta}\in\mathbb{S}^{p}_{++}$ given the empirical covariance matrix $\bm{S}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}$. Moreover, in many practical situations the graph (thus the matrix $\bm{\Theta}$) is also sparse due to limited conditional dependencies, which leads to a sparse precision matrix estimation problem.
In this case, a very popular estimator is the Graphical Lasso (GLasso) (Banerjee et al., 2008; Friedman et al., 2008) that solves the convex optimization problem

### v3

@@ Abstract
Estimating matrices in the symmetric positive-definite (SPD) cone is of interest for many applications ranging from computer vision to graph learning. While there exist various convex optimization-based estimators, they remain limited in expressivity due to their model-based approach. The success of deep learning motivates the use of learning-based approaches to estimate SPD matrices with neural networks in a data-driven fashion. However, designing effective neural architectures for SPD learning is challenging, particularly when the task requires additional structural constraints, such as element-wise sparsity. Current approaches either do not ensure that the output meets all desired properties or lack expressivity. In this paper, we introduce SpodNet, a novel and generic learning module that guarantees SPD outputs and supports additional structural constraints. Notably, it solves the challenging task of learning jointly SPD and sparse matrices. Our experiments illustrate the versatility and relevance of SpodNet layers for such applications.
@@ 1 Introduction
The estimation of symmetric positive-definite (SPD) matrices is a major area of research, due to their crucial role in various fields such as optimal transport (Bonet et al., 2023), graph theory (Lauritzen, 1996), computer vision (Nguyen et al., 2019) or finance (Ledoit and Wolf, 2003). While various statistical estimators, i.e. model-based, have been developed for specific tasks (Ledoit and Wolf, 2004; Banerjee et al., 2008; Cai et al., 2011), recent advancements focus on applying generic learning-based approaches to estimate appropriate SPD matrices with neural networks in a data-driven fashion (Huang and Van Gool, 2017; Gao et al., 2020).
@@
Training neural networks while enforcing non-trivial structural constraints such as positive-definiteness is a difficult task. There have been many efforts in this direction in recent years, often in an ad-hoc manner and each with their own shortcomings (see Section 2 for more details). Building on the seminal work of Gregor and LeCun (2010) in sparse coding, a promising research direction involves designing neural networks architectures from the unrolling of an optimization algorithm (Chen and Pock, 2016; Monga et al., 2021; Chen et al., 2022; Shlezinger et al., 2023). In the case of SPD matrices, algorithm unrolling presents several challenges. First, algorithms operating in the SPD cone usually rely on heavy operations such as retractions (Boumal, 2023), SVD or line search (Rolfs et al., 2012). These operations do not integrate well into a neural network architecture, making these algorithms difficult to unroll.
Additionally, and more importantly, many applications require further structural constraints on the learned matrix. Elementwise sparsity is a typical example of such constraints (Banerjee et al., 2008), which significantly increases the complexity of the task: learning functions that simultaneously enforce SPDness and sparsity of the output is known to be challenging (Guillot and Rajaratnam, 2015; Sivalingam, 2015). There are currently no neural architectures that enable learning jointly SPD and sparse matrices.
In this paper, we bridge this gap and make the following contributions:
  • We introduce a new SPD-to-SPD neural network architecture that also supports enforcing additional constraints on the output. We refer to it as SpodNet for Schur’s Positive-Definite Network. As a particular case, we show that SpodNet is able to learn jointly SPD and sparse matrices. To the best of our knowledge, SpodNet is the first architecture to provide strict guarantees for both properties.
@@
  • We demonstrate the framework’s relevance through applications in sparse precision matrix estimation. We highlight the limitations of other learning-based approaches and show how SpodNet addresses these issues. Our experiments validate SpodNet’s effectiveness in jointly learning SPD-to-SPD and sparsity-inducing functions, yielding competitive results across various performance metrics.
@@
We reserve the bold uppercase for matrices $\bm{\Theta}$, bold lowercase $\bm{\theta}$ for vectors and standard lowercase $\theta$ for scalars. The soft-thresholding function is $\operatorname{ST}_{\gamma}(\cdot)=\operatorname{sign}(\cdot)\max(\lvert\cdot\rvert-\gamma,0)$ for $\gamma\geq 0$; it acts elementwise on vectors or matrices. On matrices, $\lVert\cdot\rVert_{1}$ is the sum of absolute values of the matrix coefficients. The cone of $p$ by $p$ SPD matrices is denoted $\mathbb{S}^{p}_{++}$.
Figure 1: A SpodNet layer chains $p$ updates of column-row pairs and diagonals using neural networks. The matrices remain SPD at all times via Schur’s condition.
@@ 2 Related works
We first review existing approaches for estimating SPD matrices, which can be broadly divided into three categories, all suffering the same limitation of not being able to handle additional structural constraints.
Riemannian approaches
Riemannian optimization provides tools to build algorithms whose iterates lie on Riemannian manifolds, such as the SPD manifold (Absil et al., 2008). Many have adopted these tools to design neural architectures that operate on the manifold of SPD matrices (Huang and Van Gool, 2017; Gao et al., 2020; 2022). However, a common and significant bottleneck of these methods is the use of Riemannian operators, which are notoriously expensive to compute. Beyond this computational hindrance, and more importantly, current Riemannian methods are not able to impose additional sparsity on the learned matrices without breaking the SPD guarantee.
SPD layers
Neural layers with SPD outputs have also been proposed in Dong et al. (2017); Nguyen et al. (2019). A first line of work rely on the linear mapping $\bm{X}_{k+1}=\bm{W}_{k}\bm{X}_{k}\bm{W}_{k}^{\top}$ (with the layer’s weights $\bm{W}_{k}$ having full row-rank), while some others rely on clipping the eigenvalues of their output. Unfortunately, there are no obvious ways to incorporate additional structure such as elementwise sparsity on the matrices without risking breaking their SPD guarantee. For instance, hard-thresholding the off-diagonal entries of a SPD matrix does not in general preserve the SPD property (Guillot and Rajaratnam, 2012) and, more generally, elementwise functions preserving this property are very limited (Guillot and Rajaratnam, 2015).
Unrolled neural architectures
In order to ensure that a network’s output strictly respects some desired properties, a successful direction is to unroll convex optimization algorithms (Chen et al., 2022; Shlezinger et al., 2023). This unrolling procedure acts as an “inductive bias” on the architecture, naturally forcing the model to explore suitable solutions within the space of imposed properties. In SPD learning, this approach has been exploited by Shrivastava et al. (2020) who unrolled an optimization algorithm to train neural networks to estimate inverses of covariance matrices. Whilst algorithm unrolling is a very powerful approach to learn specific matrices, it proves difficult to actually enforce several constraints simultaneously on these matrices. We provide further details about the limits of this approach in Section 4.1.
@@ 3 The SpodNet framework
We now introduce our core contribution, the SpodNet layer. Essentially, it is an SPD-to-SPD mapping parameterized by neural networks. A SpodNet layer operates by cycling through the $p$ column-row pairs individually by
  1. (i)
    updating the corresponding column-row with a neural network,
  2. (ii)
    updating the diagonal element to satisfy the SPD constraint (see Figure 1).
Crucially, the neural network used at step (i) can update the column-row to any value without compromising the SPD guarantee. Consequently, one is free to exploit a spectrum of approaches for those updates, depending on other desired structural properties of the output matrix. In particular, we describe in Section 4 three specific implementations of SpodNet that enforce elementwise sparsity for learning sparse precision matrices.
@@
Algorithm 1 The SpodNet layer
@@ 3.1 Algorithmic foundations
The key to preserving the positive-definiteness $\bm{\Theta}\in\mathbb{S}_{++}^{p}$ of the matrix upon after changing its $i$-th column and row is an appropriate update of the diagonal entry $\Theta_{ii}$ based on Schur’s condition for positive-definiteness. In the following, a $+$ superscript on a variable (e.g. $\bm{\theta}^{+}$) indicates an update value of this variable (e.g. outputted by a neural network). The following proposition shows that SpodNet’s output is guaranteed to be SPD.
@@
(for a generic column $i$, $\bm{\Theta}_{11}$ refers to $\bm{\Theta}$ without its $i$-th row and $i$-th column, $\bm{\theta}_{12}$ is the $i$-th row of $\bm{\Theta}$ without its $i$-th value, and $\theta_{22}$ is $\Theta_{ii}$ as illustrated below Algorithm 1).
Suppose that $\bm{\Theta}\in\mathbb{S}_{++}^{p}$. Let $\bm{u}\in\mathbb{R}^{p-1}$ and $v>0$ be any vector and strictly positive scalar respectively. Then, updating the $i$-th row and column of $\bm{\Theta}$ as
$\bm{\Theta}^{+}\triangleq\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}^{+}_{12}\triangleq\bm{u}\\ \bm{\theta}_{21}^{+}\triangleq\bm{u}^{\top}&\theta_{22}^{+}\triangleq v+\bm{u}^{\top}[{\bm{\Theta}}_{11}]^{-1}\bm{u}\end{bmatrix}\,,$ (2)
@@
Since positivity of $\bm{\Theta}^{+}$ is guaranteed for any choice of $\bm{u}$ and $v$ as long as $v>0$, the principle of SpodNet updates is to learn these quantities as function of the current iterate, i.e. from the value of $\bm{\Theta}$, $\bm{\Theta}_{11}$, $\bm{\theta}_{12}$, etc (explicit forms for each variant will be given in Section 4.2). For simplicity, though they can depend on additional information, we write $\bm{u}=f(\bm{\Theta})$ and $v=g(\bm{\Theta})$, with $f$ and $g$ being learned mappings.
@@
Since $\bm{\Theta}$ is symmetric positive-definite, so is its leading principal submatrix $\bm{\Theta}_{11}$. It follows that $\bm{\Theta}^{+}$ is well-defined, and obviously symmetric. Its positive-definiteness ensues from Schur’s condition for positive-definiteness. Indeed $\bm{\Theta^{+}}$ is SPD when $\bm{\Theta}_{11}\succ 0$ and $\theta_{22}^{+}-\bm{\theta}_{12}^{+\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}>0$ (Zhang, 2006, Theorem 1.12). The latter condition is ensured as the left hand side is equal to $v$, which is strictly positive by assumption. ∎
Finally, one SpodNet layer chains $p$ updates of the form Equation 2, sequentially updating all column-row pairs one after the other as summarized in Algorithm 1. A full SpodNet architecture then stacks $K$ SpodNet layers, and, in practice, the first layer takes as input $\bm{\Theta}_{\mathrm{in}}=\left(\bm{S}+\mathbf{I}_{p}\right)^{-1}$ where $\bm{S}$ is the empirical covariance matrix. The overall architecture is schematized in Figure 1 for $K=2$.
The expressivity of SpodNet comes from the flexibility to use any arbitrary functions $f$ and $g$ in the updates whilst guaranteeing that $\bm{\Theta}$ remains in the SPD cone. Namely, additional structure such as sparsity can trivially be imposed on $\bm{\Theta}$ through $f$. We next detail the computational cost of fully updating $\bm{\Theta}$ through one SpodNet layer.
3.2 Improving the update complexity
The diagonal update of Equation 2 a priori requires inverting the matrix $\bm{\Theta}_{11}$ which comes with a prohibitive cost of $\mathcal{O}(p^{3})$ for each column-row update. Fortunately, we are able to leverage the column-row structure of the updates to decrease the cost to a mere $\mathcal{O}(p^{2})$, by maintaining the matrix $\bm{W}=\bm{\Theta}^{-1}$ up-to-date along the iterations.
@@ 3.2 Improving the update complexity
Let $\bm{W}$ be the inverse of $\bm{\Theta}$, adopting the same block structure $\bm{W}=\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{12}^{\top}&w_{22}\end{bmatrix}$. Then $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$. In addition, if $\bm{\Theta}$ is updated as $\bm{\Theta}^{+}$ following Equation 2 with $v=g(\bm{\Theta})$, then the update of $\bm{W}$ defined by
$\bm{W}^{+}\triangleq\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}\bm{\theta}_{21}^{+}[{\bm{\Theta}}_{11}]^{-1}}{g(\bm{\Theta})}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}}{g(\bm{\Theta})}\\ \left(-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}}{g(\bm{\Theta})}\right)^{\top}&1/g(\bm{\Theta})\end{bmatrix}\,.$ (3)
@@
By the Banachiewicz inversion formula on Schur’s complement (Zhang, 2006, Thm 1.2),
$\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{21}&w_{22}\end{bmatrix}$ $=\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix}^{-1}=\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\\ \\ \left(-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\right)^{\top}&\frac{1}{\theta_{22}-\bm{{\theta}}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\end{bmatrix}\,.$ (4)
Identifying all blocks yields $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$ which can be computed in $\mathcal{O}(p^{2})$ if one has access to $\bm{W}$. This can be achieved using Equation 3, and involves only operations in $\mathcal{O}(p^{2})$. The property $\bm{\Theta}^{+}=[\bm{W}^{+}]^{-1}$ is satisfied by using the same inversion formula. ∎
A full update of $\bm{\Theta}$ can thus be achieved with cost $\mathcal{O}(p^{3})$ ($p$ column-row updates of cost $\mathcal{O}(p^{2})$).
4 Using SpodNet to learn sparse precision matrices
So far, we have not specified which choices of column and diagonal updates $f(\bm{\Theta})$ and $g(\bm{\Theta})$ we would use in practice. We now leverage the general framework of the SpodNet layer for learning SPD matrices with additional structure and propose three specific architectures for learning sparse and SPD matrices, with applications to sparse precision matrix learning.
4.1 Inferring sparse precision matrices
Consider a dataset of $n$ observed signals $\bm{x}_{1},\ldots,\bm{x}_{n}$ where each $\bm{x}_{i}\in\mathbb{R}^{p}$ follows a certain (centered) distribution with covariance matrix $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$. The problem of precision matrix estimation arises when attempting to identify the conditional dependency graph of the $p$ variables of this dataset. In Gaussian graphical models (Lauritzen, 1996), this graph is associated with the so-called precision matrix $\bm{\Theta}=\bm{\Sigma}^{-1}\in\mathbb{S}^{p}_{++}$. In this case $\Theta_{ij}=0$ iff the variables $i$ and $j$ are conditionally independent given the other variables. This estimation problem involves determining $\bm{\Theta}$ given the empirical covariance matrix $\bm{S}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}$. In most practical scenarios, the matrix $\bm{\Theta}$ is sparse due to limited conditional dependencies, which leads to a sparse SPD matrix estimation problem. For this problem, a very popular estimator is the Graphical Lasso (GLasso) (Banerjee et al., 2008; Friedman et al., 2008) that solves